Find the fundamental, first overtone and second overtone frequencies of an open organ pipe of length 20 cm. Speed of sound in air is 340 m s-1.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. Here given distance between two nodes is = 4.0, ⇒ λ = 2 * 4.0 = 8 cm We know that v = nλ ⇒ η = 328/8 * 10^-2